Calculate the total electric potential at the center of:

a. A line segment of length 1.0 m with a +1-C charge and a –1-C charge at its endpoints; 

b. An equilateral triangle of side length 1.0 m with alternating +1-C, –1-C, and +1-C charges 

placed at its vertices;

c. A square of side length 1.0 m with alternating +1-C, –1-C, +1-C, and –1-C charges placed 

at its vertices

d. How did your answers in a, b, and c compare with your electric field calculations for similar 

configurations?


1
Expert's answer
2021-06-14T15:03:43-0400

Gives

(a) "r_1=r_2=0.5m"

"V_1=\frac{kq}{r_1}"

Put value

"V_1=\frac{9\times10^9\times 1}{0.5}=18\times10^{9}V"

"V_2=\frac{kq}{r_2}"


"V_1=-\frac{9\times10^9\times1}{0.5}=-18\times10^{9}V"

Net electric potential at symmet of center

"V_{net}=V_1+V_2"

Put value


"V_{net}=18\times10^{9}-18\times10^{9}=0V"

(b)

Centre at Point P

Point AputQ1= +1C

Side(a)=1m

Distance BP=CP=AP="\frac{1}{\sqrt3} m"

Due to +1C centre of potential

"V_A=\frac{kQ}{r_{AP}}"

"V_A=\frac{9\times10^9 \times1}{\frac{1}{\sqrt3}}=9\sqrt3V"

Similarly point B put charge Q2=-1C

"V_B=-\frac{9\times10^9 \times1}{\frac{1}{\sqrt3}}=-9\sqrt3V"

Point c put +1C

"V_c=\frac{9\times10^9 \times1}{\frac{1}{\sqrt3}}=9\sqrt3V"

"V_{net}=V_A+V_B+V_C"


"V_{net}=9\sqrt3-9\sqrt3+9\sqrt3=9\sqrt3 V"

(C)

PointA(+1C) point B(-1C)

PointC(+1C) point D(-1C)

Centre point P

AP=BP=CP=DP ="\frac{1}{\sqrt2}" m

"V_A=\frac{KQ}{r_{AP}}"

Put value

"V_A=\frac{9\times10^9\times1}{\frac{1}{\sqrt2}}=9\sqrt2V"

Similarly point B

"V_B=-\frac{9\times10^9\times1}{\frac{1}{\sqrt2}}=-9\sqrt2V"

Similarly point C

"V_C=\frac{9\times10^9\times1}{\frac{1}{\sqrt2}}=9\sqrt2V"

Similarly point D

"V_D=-\frac{9\times10^9\times1}{\frac{1}{\sqrt2}}=-9\sqrt2V"

Net potential at centre

"V=V_A+V_B+V_C+V_D"

V=

"9\sqrt2V-9\sqrt2V+9\sqrt2V-9\sqrt2V=0V"

Part(d)

Electric field

"E_A=\frac{kq}{r^2}"

Put value

"E_A=\frac{9\times10^9\times1}{.5^2}=3.6\times10^{10}N/C"


"E_B=-\frac{9\times10^9\times1}{.5^2}=-3.6\times10^{10}N/C"

"E_{net}=E_A-E_B=0"

Electric field of triangle

"E_A=\frac{kQ}{r^2_{AP}}"

"E_A=\frac{9\times10^9\times1}{(\frac{1}{\sqrt3})^2}=27" N/C"E_B=-\frac{9\times10^9\times1}{(\frac{1}{\sqrt3})^2}=-27N/C"

"E_c=\frac{9\times10^9\times1}{(\frac{1}{\sqrt3})^2}=27N/C"

"E_{net}=E_A+E_B+E_c"

Put value

"E_{net}=27N\/C"

Electric field square

"E_A=\frac{KQ}{r^2_{AP}}"

"E_A=\frac{9\times10^9\times1}{(\frac{1}{\sqrt2})^2}=18N/C"

"E_B=-\frac{9\times10^9\times1}{(\frac{1}{\sqrt2})^2}=-18N/C"

"E_C=\frac{9\times10^9\times1}{(\frac{1}{\sqrt2})^2}=18N/C"

"E_D=-\frac{9\times10^9\times1}{(\frac{1}{\sqrt2})^2}=-18N/C"

"E_{net}=E_A+E_B+E_C+E_D"

Put value

"E_{net}=0N\/c"